A mass of 7 kg is placed on an incline with friction and no additional forces acting on it and it begins to slide at some angle when the angle is slowly increased as discussed in class. Now, the same mass is placed on the same incline, but an additional force of 28 Newtons is applied to it perpendicular from the incline, away from the incline (its always in that direction). As the angle of the incline is slowly increased, it slides at a different angle with this force. If the angle which it begins to slide with the force is 35 degrees, how much further can the angle be raised with the force than without the force? Answer in degrees
Answer #1

let's figure out the equation in the first case,

mgsin \theta =\mu mg cos \theta, \mu = tan\mu

The second case,

mgcos \theta = N + 28

N =  mgcos \theta - 28

fs = \mu ( mgcos \theta - 28 )

mg sin \theta = \mu ( mgcos \theta - 28 ) -------------for the mass to slide

7 (9,8) sin 35 = \mu ( 7 x 9,8 x cos 35 - 28)

39.347 = \mu ( 28.193)

\mu= 1.4 apprx

angle in first case = tan^-1 ( 1.4) =54.438 degree apprx

your question appears a bit ambiguous, however, i have solved the parts which seemed logical to me,probbaly, the question is looking for the difference in two angles. which should be (54.438 - 35)